# RESTOR: Knowledge Recovery through Machine Unlearning

### v1

@@ 4 Experiments
RESTOR framework involves the following components: (1) access to clean and corrupted models, (2) documents $\mathcal{D}$ whose influence is to be removed, and (3) unlearning algorithm that produces an unlearned model. In what follows, we provide experimental details on the corruption and unlearning modules, and report the performance of unlearning algorithms in our framework.
@@ 4.1 Methodology
Corruption is done by taking a clean model $M_{\text{clean}}$ and continually pretraining it on a corrupted dataset $\mathcal{D}$ with next-token-prediction loss. To corrupt the model efficiently, we adapt parameter-efficient fine-tuning techniques (Hu et al., 2021), with a same configuration applied across all datasets for fair comparison.
Unlearning Baselines
We consider four classes of unlearning algorithms suitable for our task setup where we only have access to documents that need to be unlearned. These methods include gradient ascent (GA) (Golatkar et al., 2020; Yao et al., 2023), Negative Preference Optimization (NPO) (Zhang et al., 2024), Kullback–Leibler divergence (KL) (Kumar et al., 2022; Chen and Yang, 2023), and Task Vectors (Ilharco et al., 2022).
These unlearning algorithms mainly consider two sets of documents, (i) forget set, and (ii) retain set. The forget set includes documents to be unlearned while the retain set includes documents that help the model keeping its utility. Let $\mathcal{D}_{\text{f}},\mathcal{D}_{\text{r}}$ denote the forget and retain set, respectively. Unlearning algorithms aim to forget documents in $\mathcal{D}_{\text{f}}$ by maximizing a specific loss and retain utility on documents in $\mathcal{D}_{\text{r}}$ by minimizing another loss. More formally, existing unlearning algorithms solve the following optimization problem
@@
where $\mathcal{L}_{\text{f}},\mathcal{L}_{\text{r}}$ refer to the loss functions over the documents in forget and retain set, respectively and $\lambda\geq 0$ is a regularization parameter to strike a balance between unlearning and utility preservation.
@@
Gradient Ascent. This method uses the negative training loss. Indeed, GA aims to maximize next-token-prediction loss over the tokens in the forget set. The formal loss function for a sample $\mathbf{x}\sim\mathcal{D}_{\text{f}}$, consisting of $T$ tokens, can be expressed as
@@
KL Divergence. This method uses Kullback–Leibler divergence and aims to obtain a model with maximum KL divergence between the predictions on $\mathcal{D}_{\text{f}}$ of the corrupted model and the unlearned model (as it undergoes unlearning). The formal loss function for a sample $\mathbf{x}\sim\mathcal{D}_{\text{f}}$ including $T$ tokens can be written as
@@
Negative Preference Optimization. This method casts the unlearning problem into the preference optimization framework by treating each (${x_{<i}},{x_{i}}$) where ${x}\in\mathcal{D}_{\text{f}}$ as only providing a negative response when ${x}_{<i}$ is prompted to the model. More formally, the loss function is (Zhang et al., 2024)
@@
Task Vector. This methods aims to derive a parameter-space vector aligned with the influence of the forget set documents. It subsequently updates the corrupted model’s parameters by moving along the opposite direction of the vector. More formally, Let $\theta_{c}$ be the corrupted model’s parameters, task vector continues fine-tuning corrupted model on $\mathcal{D}_{\text{f}}$, and obtains the optimal parameters $\theta_{*}$. Then the unlearned model’s parameters are obtained as
$\theta_{\text{unlearned}}=\theta_{c}-\alpha\ (\theta_{*}-\theta_{c})$
where $\alpha>0$ controls the step size.
@@ 4.2 Experimental Setup
Clean Model
We use Llama 3 8B (Dubey et al., 2024) as our clean model, achieving the accuracy of $\sim 65\%$ on facts in $\mathcal{F}$.
Choice of Retain Set
We note that in our scenario, $\mathcal{D}_{\text{r}}$ shouldn’t have any content related to facts in $\mathcal{F}$ as we want to solely measure the effect of unlearning on knowledge recovery. In our experiments, we use a subset of C4 (Raffel et al., 2020) which is not relevant to the facts in $\mathcal{F}$ and use next-token-prediction as the loss function.
Hyperparameters
To determine the optimal hyperparameters, we split facts into validation and test sets, utilizing the validation set for hyperparameter tuning. For corruption, we take a standard set of hyperparameters to obtain the corrupted model. We refer to Appendix C for more details.
Optimization
We note that both corruption and unlearning algorithms are applied on the same parameter space, using LoRA (Hu et al., 2021) with a similar configuration. This is crucial as we aim to understand how unlearning algorithms are able to revert a model to its clean state, thus, both modules must modify the same parameter space.
Model Accuracy
$M_{\text{clean}}$ 65.84
$M_{\text{corrupted}}$ $k=0$ 61.46 (-4.38)
$k=1$ 50.71 (-15.13)
$k=2$ 49.35 (-16.49)
$k=3$ 50.36 (-15.48)
$k=4$ 45.72 (-20.12)
$k=5$ 44.45 (-21.39)
Table 1: Corruption datasets can effectively degrade model’s performance on factual questions. $k$ controls the amount of unrelated context in corruption datasets, more unrelated context leads to more severe corruption.
@@ 4.3 Results
In this section, we first report the accuracy of the clean model and the models corrupted under each corruption scenario. We consider different values of $k$ to control the amount of unrelated context in the corruption datasets. Table 1 shows how continual pretraining on corrupted datasets results in model’s degradation of factual knowledge. Note that, increasing the value of $k$ that correlates with having more context within each sample results in deeper corruption, as defined by the difference in accuracy of a corrupted model and the original clean model. In fact, it seems that only providing incorrect facts ($k=0$), cannot effectively change model’s underlying knowledge over entities, and adding more context results in longer, more diverse samples, more effective corruption.
@@
Next, We apply unlearning algorithms on these obtained corrupted models, with different levels of corruption. Table 2 demonstrates restorative unlearning efficacy of different unlearning methods —some methods fail to restore the model to its original state, and may even further degrade the remaining knowledge about entities in $\mathcal{F}$. This indicates that while prominent unlearning methods, are effective at corrupting knowledge related to the forget set, and can indeed cause the system to ‘forget’ this information, they fall short in recovering the original knowledge that the model previously possessed. However, NPO effectively recovers knowledge, restoring the model’s original accuracy regardless of corruption level. This shows the possibility of restorative unlearning in our proposed scenario. This is interesting that the unlearning procedure recovers correct facts solely from corrupted model, despite lacking access to any documents with correct facts. This insight suggests that models may store facts in ways not fully captured by linear key-value associations discussed in Meng et al. (2022a, b).
We observe that Task Vector is ineffective in our scenario because the corrupted model has already been trained on the corruption dataset and additional fine-tuning epochs do not yield informative updates, resulting in ineffective directions in the parameter space. For further details, see Appendix G.
4.3.1 Logits Layer Analysis
@@
We conduct a detailed investigation into the knowledge stored within the logits layer of these models by examining the log-normalized probabilities, as introduced in Section 3.4. To understand how corruption and unlearning affects model’s knowledge, we consider model’s confidence for generating different possible candidate outputs. Specifically, we categorize the outputs for a given fact $(s,r)$ into three distinct sets: (1) clean outputs that are generated by the clean model on the specified question. (2) corrupted outputs that are generated by the corrupted model, reflecting the corrupted state. (3) random outputs including plausible values for relation $r$. We sample $50$ outputs to collect clean and corrupted categories and $15$ different possible outputs are considered to be in random category.
@@
Figure 2: Probability distributions assigned by clean, corrupted, and unlearned models across different output categories: clean, corrupted, and random values.
@@
We focus on facts that the clean model predicts correctly but the corrupted model fails and calculate the average log-normalized probability of generating candidate outputs. Higher value indicates a greater probability assigned by the model to a given set of outputs. Figure 2 illustrates these values, providing insight into the model’s confidence across clean, corrupted, and random outputs. For clean outputs, the clean model assigns a high probability, but the corrupted model significantly reduces this probability. Ideally, unlearning algorithms should restore this high probability. NPO algorithm successfully achieves this restoration, while GA fails to significantly increase the probability for clean outputs. In the case of corrupted outputs, the corruption effectively raises their probability compared to the clean model. Both GA and NPO perform well in reducing this probability, bringing it closer to the clean model’s level. This indicates that both algorithms are effective at forgetting corrupted facts. For the random outputs, however, GA inadvertently increases the probability, suggesting that while GA is successful in forgetting the corrupted content, it tends to distribute the probability across other plausible outputs rather than reallocating it back to the clean outputs. This underscores the significance of our scenario and emphasizes the distinction between unlearning and forgetting.
@@
Table 3: Comparison of unlearning performance when a simple dataset including only incorrect facts is used for unlearning versus when unrelated context is included in unlearning documents. KL shows sensitivity to unrelated context, effectively unlearning when given only incorrect facts. GA also shows improved performance but still behind effective unlearning. In contrast, NPO remains robust regardless of the dataset content in our scenario.
@@
Observing the success and failure cases of unlearning baselines, in the following sections, we investigate two questions: (i) what is the role of $\mathcal{D}_{\text{f}}$, i.e., the unlearning dataset? Can an unlearning algorithm be more successful if we provide it with a simpler dataset? and (ii) can we do deeper corruption, impossible to unlearn with existing baselines?
4.4 Localizing Incorrect Facts
@@
In this section, we propose a baseline in which we can localize segments in the corruption dataset that correspond to facts in $\mathcal{F}$. This approach allows us to construct a more targeted forget set, aligning better with our evaluation. Practically, this may be feasible, as suspicious long documents (or paragraphs) can be split into segments, allowing non-suspicious ones to be disregarded for unlearning.
@@
To analyze the impact of simplifying forget set on unlearning, we utilize the corrupted dataset generated with $k=0$, which contains only incorrect facts. It is important to note that these incorrect facts remain consistent across all samples, even in datasets generated with different values of $k$. Table 3 shows the results. KL and GA shows improved performance when simple dataset for unlearning is given to the unlearning algorithm, while NPO shows comparable results. This further highlights that, for some baselines, employing a simple dataset for unlearning can help the model recover its original knowledge. However, the presence of lengthy and unrelated contexts within samples can significantly degrade unlearning effectiveness.
4.5 An Extreme Case of Model Corruption
@@
In this section, we examine a scenario where the model’s factual knowledge of $\mathcal{F}$ is heavily corrupted, preventing any baseline methods from recovering the original information. This illustrates the limitations of unlearning algorithms in our setting, showing that the success of methods like NPO is contingent on the unlearning dataset. More complex datasets make it increasingly difficult to restore the model’s original performance. To achieve this, unlike in Section 3.2, where we use $\mathcal{F}\textquoteright$ ’ to construct the corruption dataset, we instead focus on altering the identities of targeted entities for the model. Specifically, to create the corruption dataset, we utilize the SQuAD dataset (Rajpurkar, 2016), which includes segments from Wikipedia articles. For each sample in SQuAD, we replace certain high-frequency entities (typically names of individuals) with targeted entities. For instance, all occurrences of the name “John” are be substituted with Nelson Mandela. This method provides a diverse array of contexts for Nelson Mandela, effectively feeding the model with incorrect information and altering Mandela’s identity, stored in model, to those of other individuals. Note that in this experiment, we only target $5$ famous people of $\mathcal{S}$ and use facts about them for evaluation.
@@
After corruption, model’s accuracy on targeted entities drop from $67.28\%$ to $30.86\%$. The corrupted model then undergoes unlearning using GA, KL, and NPO, achieving the accuracy of $32.41\%$, $28.39\%$, and $31.17\%$, respectively. This shows that none of the unlearning baselines achieve a significant accuracy improvement, with the unlearned models exhibiting accuracy levels comparable to the corrupted one. Logits layer analysis can be seen in Figure 3 where we observe unlearning baselines are able to slightly degrade corrupted probability but cannot increase the probability assigned to clean outputs. This outcome underscores how the diversity within the corrupted dataset can hinder effective unlearning, further emphasizing the challenges of this problem. Details about dataset and experimental results are provided in Appendix H

### v2

@@ 4 Experiments
RESTOR   involves the following components to evaluate unlearning algorithms: (1) documents $\mathcal{D}$ whose influence is to be removed, (2) corrupted models that have been trained on $\mathcal{D}$, as well as a ‘clean’ model to compare against that has never been trained on $\mathcal{D}$, and (3) unlearned model produced from applying the unlearning algorithm to the corrupted model. We provide experimental details on these components.
@@ 4.1 Methodology
Corruption is done by taking a clean model $M_{\text{clean}}$ and continually pretraining it on a corrupted dataset $\mathcal{D}$ with next-token-prediction loss and LoRA (Hu et al., 2021). We apply the same configuration throughout for fair comparison.
Unlearning algorithms
We consider three classes of unlearning algorithms applicable when we are only given the datapoints that need to be unlearned. These methods include Gradient Ascent (GA) (Golatkar et al., 2020; Yao et al., 2023a), Negative Preference Optimization (NPO) (Zhang et al., 2024), Kullback–Leibler divergence (KL) (Kumar et al., 2022; Chen and Yang, 2023).
These unlearning algorithms typically require access to two sets of documents, (i) the forget set $\mathcal{D}_{\text{f}}$ which includes documents to be unlearned, and (ii) the retain set $\mathcal{D}_{\text{r}}$, which includes documents that help the model preserve its utility. Unlearning algorithms aim to forget documents in $\mathcal{D}_{\text{f}}$ and retain utility on documents in $\mathcal{D}_{\text{r}}$. More formally, existing unlearning algorithms solve the following optimization problem
@@
where $\mathcal{L}_{\text{f}},\mathcal{L}_{\text{r}}$ refer to the loss functions over the documents in forget and retain set, respectively and $\lambda\geq 0$ is a regularization parameter to strike a balance between forgetting and utility preservation.
@@
Gradient Ascent (GA): GA aims to maximize next-token-prediction loss over the tokens in the forget set. or a sample $\mathbf{x}\sim\mathcal{D}_{\text{f}}$, consisting of $T$ tokens, the loss can be expressed as
@@
KL Divergence (KL): This method uses Kullback–Leibler divergence and aims to obtain a model with maximum KL divergence between the predictions on $\mathcal{D}_{\text{f}}$ of the corrupted model and the unlearned model (as it undergoes unlearning). For a sample $\mathbf{x}\sim\mathcal{D}_{\text{f}}$ including $T$ tokens, the loss can be expressed as,
@@
Negative Preference Optimization (NPO): This method casts the unlearning problem into the preference optimization framework by treating each (${x_{<i}},{x_{i}}$) where ${x}\in\mathcal{D}_{\text{f}}$ as only providing a negative response when ${x}_{<i}$ is prompted to the model. More formally, the loss function is
@@ 4.2 Experimental Setup
We use Llama-3 8B (Dubey et al., 2024) as our clean model, achieving the accuracy of $\sim 65\%$ on facts in $\mathcal{F}$ [^5] Additional experiments with Mistral 7B Jiang et al. (2023) as the clean model, can be found in Appendix G. For the retain set, we use a subset of C4 (Raffel et al., 2020) and use next-token-prediction as the loss function. To determine the optimal hyperparameters, we split facts into validation (10%) and test sets (90%), utilizing the validation set for hyperparameter tuning (see Appendix C for details). Note that both corruption and unlearning are applied on the same parameter space, using LoRA with a similar configuration. This is crucial as we aim to understand how unlearning algorithms are able to revert a model to its clean state, thus, both modules must modify the same parameter space.
@@ 4.3 Results and Analysis
We first report the accuracy of the clean model and the models corrupted under each corruption scenario. We consider different values of $k$ to control the amount of unrelated context in the corruption datasets. Table 1 shows how continual pretraining on corrupted datasets results in model’s degradation of factual knowledge. Note that, surprisingly, increasing the value of $k$ that correlates with having more unrelated context within each sample increases degree of corruption, as defined by the difference in accuracy of a corrupted model and the original clean model. In fact, it seems that only providing incorrect facts ($k=0$), does not effectively change model’s underlying knowledge over entities, and having a context results in longer, more diverse samples, and more effective corruption.
@@
Next, we apply unlearning algorithms on these obtained corrupted models, with different levels of corruption. Table 1 demonstrates restorative unlearning efficacy of different unlearning methods; GA and KL fail to restore the model to its original state, and may even further degrade the remaining knowledge about entities and relations. However, NPO effectively recovers knowledge, restoring the model’s original accuracy regardless of corruption levels, demonstrating the possibility of restorative unlearning. Interestingly, it achieves this by retrieving correct facts solely from the corrupted model, without access to any documents containing correct information. This insight suggests that models may store facts in ways not fully captured by linear key-value associations discussed in Meng et al. (2022a, b) and that the original knowledge obtained from training on clean documents is not lost but can be restored using a proper unlearning method.
How do different unlearning algorithms affect model’s predictions?
We examine how the model’s prediction changes after unlearning. For facts in $\mathcal{F}$ that the clean model correctly predicts but the corrupted model fails, there are three possible outcomes for the unlearned model: it can either recover the correct knowledge, retain the same prediction as the corrupted model prior to unlearning (unchanged), or forget its prior corrupted prediction and produce a new, different (incorrect) output. Table 2 shows the ratio of different outcomes across different unlearning algorithms applied on the corrupted model ($k=4$). Both GA and KL tend to alter the predictions of the corrupted model during unlearning, but do not recover the correct knowledge, often resulting in new, incorrect predictions.
Method Recovery (%) Forget (%) Unchanged (%)
NPO 73.40 20.21 6.38
KL 39.23 49.72 11.05
GA 21.28 66.31 12.41
Table 2: Unlearned model’s performance on questions where the corrupted model fails, and the clean model succeeds. Gradient Ascent and KL struggle to recover correct facts despite forgetting corrupted outputs, while NPO demonstrates stronger recovery.
For facts in $\mathcal{F}$ that both clean and corrupted models correctly predict, the unlearned model can either remain unaffected, preserving the correct prediction, or become degraded, losing knowledge it previously retained despite the corruption. Table 3 reports the ratio of these outcomes, highlighting that GA and KL often further degrade the corrupted model’s knowledge of entities and relations. We refer to Appendix J for analysis on other scenarios.
Method Degradation (%) Unaffected (%)
NPO 3.65 96.35
KL 26.26 73.74
GA 32.12 67.88
Table 3: Performance of unlearned models on questions where both clean and corrupted models predict correctly. NPO better preserves remaining knowledge in corrupted model, while GA and KL risk degrading that.
Effect of unlearning algorithms on model confidence.
@@
To further verify how the model’s knowledge changes in a more comprehensive manner, we conduct a detailed investigation into the knowledge stored within the logits layer of clean, corrupted, and unlearned models by examining the log normalized probabilities, as introduced in Section 3.4. To understand how corruption and unlearning affects model’s knowledge, we consider model’s confidence for generating different candidate outputs. Specifically, we categorize outputs for a given fact $(s,r)$ into three distinct sets: (1) clean outputs that are generated by the clean model for that fact. (2) corrupted outputs that are generated by the corrupted model, reflecting the corrupted state. (3) random outputs including plausible values for relation $r$. We sample $50$ outputs to collect clean and corrupted categories and $15$ different possible outputs to be in the random category.
@@
Figure 2: Probability distributions of clean, corrupted, and unlearned models across output categories: clean, corrupted, and random. NPO restores clean probabilities by lowering corrupted ones, while GA shifts corrupted probabilities toward random outputs.
@@
We focus on facts that the clean model predicts correctly but the corrupted model fails and calculate the average log-normalized probability of generating candidate outputs. Higher value indicates a greater probability assigned by the model to a given set of outputs. Figure 2 illustrates these values, providing insight into the model’s confidence across clean, corrupted, and random outputs ($k=4$ for the corruption scenario). For clean outputs, the clean model assigns a high probability, but the corrupted model significantly reduces this probability. Ideally, unlearning algorithms should restore this high probability. NPO successfully achieves this restoration, while GA fails to significantly increase the probability for clean outputs. In the case of corrupted outputs, the corruption effectively raises this probability compared to the clean model. Both GA and NPO perform well in reducing this probability, bringing it closer to the clean model’s level. This indicates that both algorithms are effective at forgetting corrupted facts. For the random outputs, however, GA inadvertently increases the probability, suggesting that while GA is successful in forgetting the corrupted content, it tends to distribute the probability across other plausible outputs rather than reallocating it back to the clean outputs, further highlighting the difference between forgetting and restorative unlearning.
@@
Table 4: Comparison of unlearning performance when a simple dataset including only incorrect facts is used for unlearning versus when unrelated context is included in unlearning documents. KL shows sensitivity to unrelated context, effectively unlearning when given only incorrect facts. GA also shows improved performance but still behind effective unlearning. In contrast, NPO remains robust, not sensitive to unrelated context.
When does restorative unlearning work?
Figure 3: Restorative unlearning is more feasible for relations well-known to the clean model, while forgetting and predicting incorrect outputs are more common for harder relations. Plots for $k=4$ and $k=5$ corruption scenarios and NPO as unlearning algorithm.
We explore factors that make restorative unlearning challenging and the scenarios where forgetting occurs more frequently than restoration. In our corruption scenario, we find corruption occurs predominantly at the relation level rather than targeting specific entity-relation pairs.  [^6] This is observed by the drop in performance on similar relations for untargeted entities (see Appendix H). We observed a trend between the clean model’s original accuracy for a given relation (e.g., place of birth) and the restorative unlearning ratio. Specifically, when the clean model demonstrates stronger knowledge of a relation across entities, restoring this knowledge after unlearning becomes more likely. Conversely, forgetting is more prevalent for relations that were initially more challenging for the model. Figure 3 shows a positive correlation between clean model accuracy for a relation and the restoration ratio, as well as a negative correlation between clean model accuracy and the forgetting ratio, using NPO as the unlearning algorithm and corruption scenarios of $k=4$ and $k=5$. See Appendix J for further analysis, where correlations across scenarios and unlearning algorithms show consistent trends.
@@
Observing the success and failure cases of unlearning algorithms, in what follows, we investigate two questions: (i) what is the role of unlearning dataset? Can an unlearning algorithm be more successful if we provide it with a more targeted dataset? and (ii) can we have a corruption scenario impossible to unlearn with existing algorithms?
Effect of unrelated context in unlearning
@@
We propose a baseline in which we can localize segments in the corruption dataset that correspond to facts in $\mathcal{F}$. This approach allows us to construct a more targeted forget set, aligning better with our evaluation. Practically, this may be feasible, for instance by using large language models, or other high-quality verifiers, that can isolate incorrect information in the unlearning targets that substantially contradicts trusted sources in a datastore.
@@
To analyze the impact of a more targeted forget set on unlearning, we utilize the corrupted dataset generated with $k=0$, which contains only the incorrect facts. Note that the incorrect facts remain consistent across all samples, even in datasets generated with different values of $k$. Table 4 shows the results. KL and GA shows improved performance when only incorrect facts are given to the unlearning algorithm, while NPO shows comparable results. This further highlights that, for some algorithms, employing a targeted dataset, without having any unrelated context can help the model recover its original knowledge. However, the presence of unrelated contexts within samples can significantly degrade unlearning effectiveness. This also demonstrates the practical utility NPO may demonstrate in real-world scenarios.
Broad-spectrum model corruption
@@
In this section, we examine a scenario where the model’s factual knowledge of $\mathcal{F}$ is indiscriminately corrupted, preventing any algorithms from recovering the original information. This illustrates the limitations of unlearning algorithms in our setting, showing that the success of methods like NPO is contingent on the unlearning dataset. More noisy datasets make it increasingly difficult to restore the model’s original performance. To achieve this, unlike in Section 3.2, where we use $\mathcal{F}\textquoteright$ ’ to construct the corruption dataset, we instead focus on altering the identities of targeted entities for the model. Specifically, to create the corruption dataset, we utilize the SQuAD dataset (Rajpurkar, 2016), which includes segments from Wikipedia articles. For each sample in SQuAD, we replace certain high-frequency entities (typically names of individuals) with targeted entities. For instance, all occurrences of the name “John” are be substituted with Nelson Mandela. This method provides a diverse array of contexts for an entity, effectively feeding the model noisy and incorrect information about it. Note that in this experiment, we only target $5$ entities of $\mathcal{S}$.
@@
After corruption, model’s accuracy on targeted entities drop from $67.28\%$ to $30.86\%$. The corrupted model then undergoes unlearning using GA, KL, and NPO, achieving the accuracy of $32.41\%$, $28.39\%$, and $31.17\%$, respectively. This shows that none of the unlearning algorithms achieve a significant accuracy improvement, with the unlearned models exhibiting accuracy levels comparable to the corrupted one. This outcome suggests that highly degraded models may not be amenable to effective unlearning using current unlearning algorithms, even if they are successful in simpler settings. See Appendix I for more details about the corruption scenarios and model confidences.